# RAIL: Robot Affordance Imagination with Large Language Models

### v1

@@ Abstract
This paper introduces an automatic affordance reasoning paradigm tailored to minimal semantic inputs, addressing the critical challenges of classifying and manipulating unseen classes of objects in household settings. Inspired by human cognitive processes, our method integrates generative language models and physics-based simulators to foster analytical thinking and creative imagination of novel affordances. Structured with a tripartite framework consisting of analysis, imagination, and evaluation, our system “analyzes” the requested affordance names into interaction-based definitions, “imagines” the virtual scenarios, and “evaluates” the object affordance. If an object is recognized as possessing the requested affordance, our method also predicts the optimal pose for such functionality, and how a potential user can interact with it. Tuned on only a few synthetic examples across 3 affordance classes, our pipeline achieves a very high success rate on affordance classification and functional pose prediction of 8 classes of novel objects, outperforming learning-based baselines. Validation through real robot manipulating experiments demonstrates the practical applicability of the imagined user interaction, showcasing the system’s ability to independently conceptualize unseen affordances and interact with new objects and scenarios in everyday settings.
@@ I INTRODUCTION
In domestic settings and healthcare facilities, unstructured environments and dynamic daily tasks necessitate human-level intelligence for robots to automatically and adaptively engage in novel objects and scenarios. Recently, Large Language Models (LLMs) have showcased their impressive conversational and logical abilities. Trained on vast amounts of data, LLMs can distill essential information from ambiguous and unspecific requests to create coherent narratives. Recent studies have demonstrated that LLMs can assist robots in making high-level decisions applicable to real-world scenarios. A key challenge lies in the fact that LLMs lack a practical grasp of physics, which hinders their comprehension of the physical world and their ability to make grounded assessments and feasible plans. For example, when provided with a description and asked “Is it a table?”, LLMs may provide a logical response that lacks physical plausibility.
@@
Integrating physical properties, the concept of robot imagination assesses the object affordances from an interactive perspective, enriching the information for robot manipulation. To approach the affordance reasoning from a user-centric perspective, we define the affordance of an object by Interaction-Based Definition (IBD). Illustrated by the examples of chairs and containers in [1, 2, 3], IBD defines the potential user and feasible user-object interactions, providing extensive instructions for the robot to imagine the scenario and assess the resultant interaction. We define the pose of the object that allows the expected interaction as the functional pose for the target affordance. The object that has at least one functional pose is recognized as functional, i.e., fulfilling such affordance.
@@
Figure 1: Robot imagination with LLMs. (a) The robot imagines the affordances of randomly placed novel objects assisted with LLMs. (b) The robot performs novel tasks based on affordance reasoning.
@@
Figure 2: Pipeline. Given an object model in a random pose, the algorithm first imagines its stable poses. The Imagination Analyzer analyzes the requested affordance and generates an executable imagination profile. The algorithm simulates the imagination profile with the object and loop for all stable poses. The Imagination Evaluation determines whether the object has the requested affordance. If the object is functional, the functional pose and agent trajectories are recorded for potential real robot execution.
@@
Current robot imagination methods require the development of customized imagination systems for different affordances, in which developers analyze the definition of a class of objects from the dictionary and decompose it into an IBD. To translate the affordance request into the programming language that the robot can understand and execute, in our previous works [2, 3], we propose the simplified agent model to describe the potential user and encode the interaction into feasible motions and expected outcomes. We define the applicable agent models and motions as the imagination profile. An evaluation matrix is tuned to determine the successful interactions, allowing the robot to recognize the functional poses of the object and, therefore, classify the objects. However, it is still an open challenge to imagine novel affordances without complicated implementations by human developers.
@@
Evaluated with 301 novel synthetic data, our method showcases a robust 88.2% accuracy in identifying new affordances and an impressive 92.7% success rate in determining functional poses. In real robot experiments, the system recognized both the affordances and the functional poses of 18 previously unseen objects. It achieves a 100% success rate in executing novel tasks by accurately parsing semantic requests and reasoning novel affordances. Comparing it with leading learning-based approaches and an ablation study baseline, we empathize the effectiveness, generality, and practical applicability of our method. The main contribution of this paper lies in:
@@
  • An imagination framework simulates customized profiles for multi-class affordances.
@@ II Related Work
II-A Learning-based Object Classification
@@
Object classification has been a crucial area of research in computer vision, a fundamental approach in this field is the application of Convolutional Neural Networks (CNNs) [4], based on which later works have made improvement and achieved better performance in tasks with large datasets [5, 6, 7, 8]. Despite the speed and efficacy of these methods, they impose strict requirements on images. One challenge is that visual occlusions can greatly impact result accuracy, and potential flaws may go unnoticed by the method, leading to objects lacking their intended functionality (e.g., a cup without a bottom cannot hold liquid). To address this, studies have employed sophisticated techniques to integrate multiple viewpoints for 3D object classification [9, 10, 11]. However, they still encounter difficulties with object poses, resulting in a significant drop in classification accuracy when objects are not in their functional poses.
Recently, the incorporation of methods such as attention mechanisms and transformer models, as demonstrated in Vision Transformer (ViT) [12], indicates a promising avenue to further enhance classification precision by utilizing global context and long-distance dependencies in images. Nevertheless, based on visual information, these methods face similar challenges. In addition, requiring substantial datasets and significant computational resources and time for training such a substantial model remain issues. In contrast, our approach does not depend on visual cues, but infers the object’s affordance through physical interaction in simulation. And by employing a pre-trained frozen LLM, our system releases the need for training process and data requirements.
II-B Affordance Recognition by Physical Reasoning
@@
The detection of affordances has recently been a highly popular subject [13, 14, 15, 16]. This area of study helps robots better identify objects in their environment, understand their functionalities, and learn to interact with them appropriately. One of the approaches to affordance detection involves physical interaction, which is seen as an intuitive and natural behavior of human [17]. Research works such as [18] and [19] suggest that robots can explore the functionalities of new objects by interacting with them, potentially using them as tools to perform tasks that were previously inaccessible. Our previous research also focuses on envisioning the affordances of objects, such as determining the sitability of a chair [2] and the containability of a cup [1], and utilizing the functional poses of these objects for robot manipulation tasks in real-world settings. While these methods are intuitive, setting up the simulation environment involves extensive hard coding and only identifies a specific affordance. In this study, we streamline the analysis and evaluation process by leveraging the reasoning capabilitiy of LLMs and applying it to robot imagination with various novel affordances.
II-C Robot Planning with Large Language Models
As interest in large language models grows, the reasoning capability of these models has provided new opportunities for robot planning, a problem that previously required complex algorithm development [20, 21, 22]. Furthermore, their language analysis skills can help robots better understand user needs based on abstract instructions [23, 24, 25]. A unified model suggested in [21] can be implemented across various robots for diverse task executions. Fan et al. highlight the potential of LLMs in industrial robot applications in [26], with the aim of autonomous completion of production tasks on different assembly lines. One work similar to ours is [27], where a tool recognition system for robot task completion is introduced that comprises multiple LLM modules including an analyzer, planner, encoder, etc. In contrast to these instances, our focus is on utilizing a framework with several LLM modules to create an autonomous system that classifies objects by imagining the affordances of novel objects with minimal human intervention.
@@ III Language Model Assisted Imagination
The pose of a rigid body can be represented as $g=(R,\mathbf{p})\in SE(3)$, where $R\in SO(3)$ is a rotation matrix, $\mathbf{p}\in\mathbb{R}^{3}$ describes the position. Functional pose $g_{f}$ is therefore defined as a pose in which the object affords the functionality. Given an unseen object and a novel affordance, our goal is three-fold: verifying whether the object possesses such affordance, identifying the functional pose, and how to use the object if it is functional.
In this work, we approach the problem of affordance reasoning from an interaction-based perspective by employing a generalized automatic robot imagination pipeline shown in Fig. 2. The novel affordance can be assessed by a three-step stream: 1) imagination analysis: analyzing the IBD of the affordance and decomposing it into the potential interaction of agents and expected outcome; 2) imagination: simulating the generated interaction in a physics-based simulator; 3) imagination evaluation: evaluating the quantitative imagination results and proposing the object affordance. Instead of heuristic reasoning from human developers, we propose a framework that integrates LLMs for analysis that can be generalized to a wide scope of affordances.
@@
In IBD, an object must be in a stable pose to be considered functional for any affordance. Therefore, the algorithm first takes the 3D model of the unseen object and finds a set of stable poses by stable pose imagination [2], referred to as $g_{s}\in G_{s}$. Setting the object into each stable pose as the candidate pose, by breaking the imagination analysis into a high-level affordance analysis and a low-level imagination profile generation, our approach proposes the four-step workflow: 1) Affordance Analyzer $\mathcal{A}(d_{\text{task}})\rightarrow{d_{\text{IBD}},d_{\text{agent}},d_{\text{interaction}},d_{\text{evl}}}$ parses the input $d_{\text{task}}$ and generates the IBD $d_{\text{IBD}}$ and an abstract imagination outline including the description of the agent $d_{\text{agent}}$ and agent action $d_{\text{interaction}}$, and expected outcome $d_{\text{evl}}$. 2) Imagination Profile Generator $\mathcal{G}(d_{\text{agent}},d_{\text{interaction}})\rightarrow{a,p}$ takes the analyzed outline as input and outputs executable agent model $a$ and trajectories $\mathcal{T}$ for imagination. 3) Imagination $\mathcal{I}(a,\mathcal{T})\rightarrow\mathcal{R}$ simulates the generated agent motions and saves resultant configurations $\mathcal{R}$. 4) Imagination Evaluator $\mathcal{E}(d_{\text{agent}},d_{\text{interaction}},d_{\text{evl}},\mathcal{R})\rightarrow{j}$ outputs the summarized affordance judgments by generating a scoring function to evaluate the imagination outcomes. The workflow is shown in Fig. 3.
@@ III-A Affordance Analysis
The Affordance Analyzer $\mathcal{A}$ processes the task description $d_{\text{task}}$ and proposes analysis as the outline of the affordance reasoning. This module first assesses whether the object dimension matches the target affordance $r_{\text{aff}}$. If there is a mismatch, it proposes a potential alternative affordance and replaces it as $r_{\text{aff}}$. Given the target affordance, the analyzer summarizes the IBD $d_{\text{IBD}}$. Based on $d_{\text{IBD}}$, it proposes a physical figure that can act as an agent to verify the target affordance. Multiple agents are involved in simulating complicated physics or efficient parallel computing. The configuration of the agent $d_{\text{agent}}$ is analyzed as a language description of the shape and the abstract layout, i.e., the geometry shape of the agent, and how the agent(s) distribute. The analyzer plans the active motion $d_{\text{interaction}}$ of the agent(s) to interact with the object and forecast the outcome, $d_{\text{result}}$. If the resultant configuration of the planned agent(s) motion aligns with the expected outcome, the object is identified as functional.
@@ III-B2 Agent Configuration Generator
To handle complicated scenarios, the imagination analyzer proposes multiple agents to simulate joint or parallel physics. We organize the spatial distribution of the agent into a cube or planar grid pattern, with sides of the shape aligning with the world frames, and the orientations of all agents remaining the same, an example is shown in Fig. 4(b). The distribution is parameterized by the number of agents along each side and the distance between each pair of neighboring agents. Based on $d_{\text{task}}$ and $a$, the LLM module proposes the distribution parameters, considering the object dimension and collision avoidance. Therefore, the pose of each agent relative to the geometric center of the distribution can be extracted. For cases where only one agent is employed, the agent geometric center is the center of the distribution.
@@ III-B3 Agent Motion Planner
Taking $d_{\text{task}}$ and the abstract description of the interaction method $d_{\text{interaction}}$, the planner aims to produce high-level plans of agent motions, shown in Fig. 4(c). Specifically, each plan provides a sequence of actions $d_{\text{motion}}$, indicating the relative spatial relations between the agent and object, as well as the moving direction of the agent. With the object placed in each candidate pose, it generates multiple plans $\mathcal{D}=\{d_{\text{motion}1},d_{\text{motion}2},...,d_{\text{motion}i},...\}$, enabling the exploration of a diverse range of interactions. $d_{\text{motion}i}$ is the language description of the i-th plan.
@@ III-B4 Trajectory Generator
With the initial language descriptions $\mathcal{D}$ of the plans, the generator converts them into a group of numerical trajectories, referred to as $\mathcal{T}=\{t_{1},t_{2},...,t_{i},...\}$. $t_{i}$ is the i-th trajectory of the center of the agent(s). Each trajectory is given as a sequence of via poses of the agent(s) relative to the object, with the object in a candidate pose. The algorithm calculates the trajectory of each agent as the executable plan saved in a JSON file imported for imagination.
@@ III-C Imagination
The algorithm imagines the proposed agent $a$ and planned interactions $\mathcal{T}$ in a physics-based simulator. It first loads the objects in the candidate stable poses and the agents in the initial poses, then simulates the agent moving through the planned via-poses and the agent is released after the trajectory is completely executed. In addition, collision checking is performed at each step, and when a collision occurs, all agents involved are released immediately.

### v2

@@ Abstract
This paper introduces an automatic affordance reasoning paradigm tailored to minimal semantic inputs, addressing the critical challenges of classifying and manipulating unseen classes of objects in household settings. Inspired by human cognitive processes, our method integrates generative language models and physics-based simulators to foster analytical thinking and creative imagination of novel affordances. Structured with a hierarchical framework consisting of analysis, imagination, and evaluation, our system “analyzes” the requested affordance names into interaction-based definitions, “imagines” the virtual scenarios, and “evaluates” the object affordance. If an object is recognized as possessing the requested affordance, our method also predicts the optimal pose for such functionality and how a potential user can interact with it. Tuned on only a few synthetic examples across 3 affordance classes, our pipeline achieves a very high success rate on affordance classification and functional pose prediction of 8 classes of novel objects, outperforming learning-based baselines. Validation through real robot manipulating experiments demonstrates the practical applicability of the imagined user interaction, showcasing the system’s ability to independently conceptualize unseen affordances and interact with new objects and scenarios in everyday settings.
@@ I Introduction
In domestic settings and healthcare facilities, unstructured environments and dynamic daily tasks necessitate human-level intelligence for robots to automatically and adaptively engage in novel objects and scenarios. Recently, Large Language Models (LLMs) have showcased their impressive conversational and logical abilities. Trained on vast amounts of data, LLMs can distill essential information from ambiguous and unspecific requests to create coherent narratives. Recent studies have demonstrated that LLMs can assist robots in making high-level decisions that are applicable to real-world scenarios. A key challenge lies in the fact that LLMs lack a practical grasp of physics, which hinders their comprehension of the physical world and their ability to make grounded assessments and feasible plans. For example, when provided with a description and asked “Is it a table?”, LLMs may provide a logical response that lacks physical plausibility.
@@
Integrating physical properties, the concept of robot imagination assesses the affordances of objects from an interactive perspective, enriching the information for robot manipulation. To approach the affordance reasoning from a user-centric perspective, we define the affordance of an object by Interaction-Based Definition (IBD). Illustrated by the examples of chairs and containers in [1, 2, 3], IBD defines the potential user and feasible user-object interactions, providing extensive instructions for the robot to imagine the scenario and assess the resultant configuration. We define the pose of the object that allows the expected interaction as the functional pose for the target affordance. The object that has at least one functional pose is recognized as functional, i.e., fulfilling such affordance.
@@
Figure 1: Overview of robot affordance imagination with LLMs. (a) The robot imagines the affordances of randomly placed novel objects assisted with LLMs. (b) The robot performs novel tasks based on affordance reasoning.
@@
Figure 2: Pipeline. Given an object model in a random pose, the algorithm first imagines its stable poses. The Imagination Analyzer analyzes the requested affordance and generates an executable imagination profile. The algorithm simulates the imagination profile with the object and loops for all stable poses. The Imagination Evaluator determines whether the object has the requested affordance. If the object is functional, the functional pose and agent trajectories are recorded for potential real robot execution.
@@
Current robot imagination methods require the development of customized imagination systems for different affordances, in which developers analyze the definition of a class of objects from the dictionary and decompose it into an IBD. To translate the affordance request into the programming language that the robot can understand and execute, in previous works [2, 3], the simplified agent model is proposed to describe the potential user and encode the interaction into feasible motions and expected outcomes. We define the applicable agent models and motions as the imagination profile. An evaluation matrix is tuned to determine the successful interactions, allowing the robot to recognize the functional poses of the object and, therefore, classify the objects. However, it is still an open challenge to imagine novel affordances without complicated implementations by human developers.
@@
Evaluated with 301 novel synthetic data, our method showcases a robust 88.2% accuracy in identifying novel affordances and an impressive 92.7% success rate in determining functional poses. In real robot experiments, the system recognizes both the affordances and the functional poses of 18 previously unseen objects. It achieves a 100% success rate in executing novel tasks by accurately parsing semantic requests and reasoning novel affordances. Comparing it with leading learning-based approaches and an ablation study baseline, we empathize effectiveness, generality, and practical applicability of our method. The main contribution of this paper lies in:
@@
  • An imagination framework that simulates customized profiles for multi-class affordances.
@@ II Related Work
II-A Object Classification
@@
Object classification is a critical area in computer vision, with Convolutional Neural Networks (CNNs) [4] forming the foundational approach. Subsequent improvements have significantly enhanced performance on large datasets [5, 6, 7, 8]. Despite their speed and efficacy, these methods impose strict requirements on images. Visual occlusions can significantly impact accuracy, and unnoticed flaws may result in objects lacking their intended functionality (e.g., a cup without a bottom cannot hold liquid). To address these issues, researchers have integrated multiple viewpoints for 3D object classification [9, 10, 11]. However, these techniques struggle with object poses, leading to reduced classification accuracy when objects are not in their functional poses.
In contrast, instead of relying on visual cues, our approach infers the object’s affordance by simulating potential physical interaction with the object in different poses. By employing a pre-trained frozen LLM, our system also releases the need for training processes and data requirements.
II-B Physical Reasoning of Object Affordance
@@
Affordance detection has recently become a popular research subject [12, 13, 14, 15, 16], aiding robots in identifying objects, understanding their functionalities, and interacting with them appropriately. One approach involves physical interaction, mirroring the intuitive behavior of humans [17]. Studies [18, 19] suggest that robots can explore new object functionalities through interaction, potentially using them as tools for previously inaccessible tasks. Previous research has focused on envisioning affordances by imagination, such as the sittability of a chair [2] and the containability of a cup [1], and utilizing these functional poses for robot manipulation tasks [3, 1]. However, developing these imagination pipelines and evaluation matrices requires extensive effort, and the systems can only identify specific affordances. This study streamlines the analysis and evaluation process by leveraging the reasoning capability of LLMs, enabling application to various novel affordances.
II-C Robot with Large Language Models
Vision-Language Models (VLMs) and LLMs have proven effective in broad robotics applications in household, industrial [20], and laboratory settings [21]. A significant area of application is language-conditioned object manipulation, where LLMs and VLMs facilitate robot manipulation by aiding in dataset preparation, task analysis, agent policy formulation, and reward function design. One major challenge in robot learning is the collection and scaling of training data. Katara et al. [22, 23] automate the generation of 3D object assets and tasks for simulation environments. LLMs define object physics parameters, meaningful tasks, and subtask details.
LLMs also serve as motion planners [24, 25, 26, 27, 28], though grounding language commands to low-level actions remains challenging. High-level task planners, such as SayCan [29], combine skill usefulness probabilities and execution success, but often require manual engineering or prior learning. Some researchers use LLMs to evaluate planning outcomes [30, 31, 32, 33], while hybrid approaches [34] leverage LLMs for task analysis and traditional planners for motion execution. Our focus is on leveraging LLMs for affordance reasoning and applying planned motions to manipulation without prior knowledge or manual engineering.
@@ III Language Model Assisted Imagination
The pose of a rigid body can be represented as $g=(R,\mathbf{p})\in SE(3)$, where $R\in SO(3)$ is a rotation matrix, $\mathbf{p}\in\mathbb{R}^{3}$ describes the position. Functional pose $g_{f}$ is therefore defined as a pose in which the object affords the functionality. Given an unseen object and a novel affordance, our goal is three-fold: verifying whether the object possesses such affordance, identifying the functional pose, and determining how to use the object if it is functional.
In this work, we approach the problem of affordance reasoning from an interaction-based perspective by employing a generalized automatic robot imagination pipeline shown in Fig. 2. The novel affordance can be assessed by a three-step stream: 1) imagination analysis: analyzing the IBD of the affordance and decomposing it into the potential interaction of agents and expected outcome; 2) imagination: simulating the generated interaction in a physics-based simulator; 3) imagination evaluation: evaluating the quantitative imagination results and proposing the object affordance. Instead of employing human developers to analyze the affordance definition, we propose an automatic framework that integrates LLMs for a wide scope of affordances.
@@
In IBD, an object must be in a stable pose to be considered functional for any affordance. Therefore, the algorithm first takes the 3D model of the unseen object and finds a set of stable poses by stable pose imagination [2], referred to as $g_{s}\in G_{s}$. Setting the object into each stable pose as the candidate pose, by dividing the imagination analysis into a high-level affordance analysis and a low-level imagination profile generation, our approach proposes the four-step workflow: 1) Affordance Analyzer $\mathcal{A}(d_{\text{task}})\rightarrow{(d_{\text{IBD}},d_{\text{agent}},d_{\text{interaction}},d_{\text{evl}})}$ parses the input $d_{\text{task}}$ and generates the IBD $d_{\text{IBD}}$ and an abstract imagination outline including the description of the agent $d_{\text{agent}}$ and agent action $d_{\text{interaction}}$, and expected outcome $d_{\text{evl}}$. 2) Imagination Profile Generator $\mathcal{G}(d_{\text{agent}},d_{\text{interaction}})\rightarrow{a,\mathcal{T}}$ takes the analyzed outline as input and outputs executable agent model $a$ and trajectories $\mathcal{T}$ for imagination. 3) Imagination $\mathcal{I}(a,\mathcal{T})\rightarrow\mathcal{R}$ simulates the generated agent motions and saves resultant configurations $\mathcal{R}$. 4) Imagination Evaluator $\mathcal{E}(d_{\text{agent}},d_{\text{interaction}},d_{\text{evl}},\mathcal{R})\rightarrow{j}$ outputs the summarized affordance judgments by generating a scoring function to evaluate the imagination outcomes. The workflow is shown in Fig. 3.
@@ III-A Affordance Analysis
The Affordance Analyzer $\mathcal{A}$ processes the task description $d_{\text{task}}$ and proposes the outline of the affordance reasoning. This module first assesses whether the object dimension matches the target affordance $r_{\text{aff}}$. If there is a mismatch, it proposes a potential alternative affordance and replaces it as $r_{\text{aff}}$. Given the target affordance, the analyzer summarizes the IBD $d_{\text{IBD}}$. Based on $d_{\text{IBD}}$, it proposes a physical figure that can act as an agent to verify the target affordance. Multiple agents are involved in simulating complicated physics or efficient parallel computing. The configuration of the agent $d_{\text{agent}}$ is analyzed as a language description of the shape and the abstract layout, i.e., the geometry shape of the agent, and how the agent(s) distribute. The analyzer plans the active motion $d_{\text{interaction}}$ of the agent(s) to interact with the object and forecast the outcome, $d_{\text{result}}$. If the resultant configuration of the planned agent(s) motion aligns with the expected outcome, the object is identified as functional.
@@ III-B2 Agent Configuration Generator
To handle complicated scenarios, the imagination analyzer proposes multiple agents to simulate joint or parallel physics. We organize the spatial distribution of the agent into a cube or planar grid pattern, with sides of the shape aligning with the world frames, and the orientations of all agents remaining the same. The distribution is parameterized by the number of agents along each side and the distance between each pair of neighboring agents. Based on $d_{\text{task}}$ and $a$, the LLM module proposes the distribution parameters, considering the object dimension and collision avoidance. Therefore, the pose of each agent relative to the geometric center of the distribution can be extracted. For cases where only one agent is employed, the agent geometric center is the center of the distribution. An example of agent configuration is shown in Fig. 4(b), where only one agent is used.
@@ III-B3 Agent Motion Planner
Taking $d_{\text{task}}$ and the abstract description of the interaction method $d_{\text{interaction}}$, the planner aims to produce high-level plans of agent motions, shown in Fig. 4(c). Specifically, each plan is given by a sequence of actions $d_{\text{motion}}$, indicating the relative spatial relations between the agent and object, as well as the moving direction of the agent. With the object placed in each candidate pose, it generates multiple plans $\mathcal{D}=\{d_{\text{motion}1},d_{\text{motion}2},...,d_{\text{motion}i},...\}$, enabling the exploration of a diverse range of interactions. $d_{\text{motion}i}$ is the language description of the i-th plan.
@@ III-B4 Trajectory Generator
With the initial language descriptions of the plans, $\mathcal{D}$, the generator converts them into a group of numerical trajectories, referred to as $\mathcal{T}=\{t_{1},t_{2},...,t_{i},...\}$. $t_{i}$ is the i-th trajectory of the center of the agent(s). Each trajectory is given as a sequence of via poses of the agent(s) relative to the object, with the object in a candidate pose. The algorithm calculates the trajectory of each agent as the executable plan saved in a JSON file, which can be imported for imagination.
@@ III-C Imagination
The algorithm imagines the proposed agent $a$ and planned interactions $\mathcal{T}$ in a physics-based simulator. It first loads the objects in the candidate stable poses and the agents in the initial poses, then simulates the agent moving through the planned via-poses. The agent is released after the trajectory is completely executed. In addition, collision checking is performed at each step, and when a collision occurs, all agents involved are released immediately.